OpenStudy (anonymous):

multiply 7/11 x 2/3=

OpenStudy (anonymous):

(a/b) * (c/d) = ac/bd

OpenStudy (anonymous):

ok when multiply do you do it straight across or do you cross multiply

OpenStudy (anonymous):

14/33

OpenStudy (anonymous):

straight across 14/33

OpenStudy (anonymous):

4 x 3 = (4/1) x (3/1) = (3 x 4)/(1 x 1) = 12/1 = 12

OpenStudy (anonymous):

thats what I have but because I couldn't simplify it I was not sure